An object moves along the x axis according to the equation x = 3, 15t2-2.00t + 3.00, where x is in meters and t is in seconds. (a) Determine the average speed between t-2.40 s and t 4.40 s m/s (b) Determine the instantaneous speed at t2.40 s m/s Determine the instantaneous speed at t4.40 s m/s (c) Determine the average acceleration between t-2.40 s and t- 4.40 s. m/s2 (d) Determine the instantaneous acceleration at t2.40 s m/s2 Determine the instantaneous acceleration at t 4.40 s. m/s2 (e) At what time is the object at rest?
Answer #1

Given is:-

The equation  

r(t) = 3.15P-2t+ 3

at t=2.40s the position of the object is

7(2.40) = 3.15( 2.40)2-2(2.40) + 3

which gives 1、(2.40) = 16.344m

similarly at t=4.40s the position is  

7(4.40) = 3.15( 4.40)2-2(4.40) + 3

which is  r(4.40) 55.184 m

thus the average speed is given by

uavg_ At

by plugging all the values we get

55.184 -16.344 4.40 2.40 aug-

avg- 19.42m/s

Part-b

V) dt

thus differentiating the given function w.r.t time we get

r(t) = 6.30t-2

now

at=2.40 the instanteous velocity will be

:2.40 6.30( 2.40)-2

(2.40)13.12m/s

similarlly

v(4.40) = 6.30(4.40)-2

U(4.40) 25.72m/s

Part-c

a = rac{Delta v}{Delta t}

25.72 - 13.12 4.40-2.40

which gives us

oxed{a = 6.3m/s^2}

Part-d

dv dt

By differentiating the velocity funtion w.r.t time we get

a(t)-6.30

thus at time t=2.40 and at t=4.40 the instantaneous acceleration will be same

therefore

a-6.30772 /s*

part-e

Object will be at rest when the velocity is zero therefore

6.30t-2-0

which gives us

| t = 0.32s